Math 331, Fall 2004

1. Compute the .

2. Write a TurtleCmd procedure that draws the n-th approximation of a fractal of your choice (not seen in class) and calculate its box-counting dimension.

3. Construct a Cantor set whose box counting dimension is 1/2. Explain a general algorithm for constructing a Cantor set with any given box dimension 0 < d < 1. [You can do this on Maple or by hand.]

4. Write an IFS procedure to draw the n-th approximation of a fractal of your choice (not one seen in class) and compute its box-counting dimension.